OpenStudy (anonymous):

medal for best answer!!!!!;):):):):) 1.Explain how you can use the distributive property to check the binomial factors when a trinomial has been factored. Include an example in your explanation

OpenStudy (campbell_st):

what you can do is split one of the binomials and then apply the distributive property if f(x) = (x + 3)(x -2) split the 1st binomial f(x) = x(x -2) + 3(x -2) now apply the distributive property to get \[f(x) = x^2 - 2x + 3x - 6\] which simplifies to \[f(x) = x^2+ x - 6\]

OpenStudy (anonymous):

thanks :) can you help me with another one????

OpenStudy (campbell_st):

just post the question in the normal manner and there are lots of people who can help